OpenStudy (anonymous):

Evaluate: \(\left[\begin{matrix}1 & 2 &1 & 3 \\2&3&2&0\\3&1&0&1\\0&0&3&2\end{matrix}\right]\)

OpenStudy (anonymous):

I got -102. Is it correct or did I get an arithmetic error somewhere?

OpenStudy (anonymous):

I got -98

OpenStudy (anonymous):

Hmm... Let me just double check everything here...

OpenStudy (anonymous):

\[3\left[\begin{matrix}1 & 2 &3\\ 2 & 3 &0\\3&1&1\end{matrix}\right] - 2\left[\begin{matrix}1 & 2 &1\\ 2 & 3 &2\\3&1&0\end{matrix}\right]\]Did you end up doing this?

OpenStudy (turingtest):

determinant?

OpenStudy (anonymous):

I think so.

OpenStudy (turingtest):

must be method of cofactors, eh?

OpenStudy (anonymous):

Yup...

OpenStudy (asnaseer):

I get 72. But maybe I made a mistake. Might be worth looking at this to help: http://people.richland.edu/james/lecture/m116/matrices/determinant.html

OpenStudy (turingtest):

I get 72 as well

OpenStudy (asnaseer):

Look down that web page at the section titled "Larger Order Determinants"

OpenStudy (turingtest):

another source for ya http://tutorial.math.lamar.edu/Classes/LinAlg/MethodOfCofactors.aspx

OpenStudy (anonymous):

Well, 72 is one of my answer choices... I did it again, but I got - 72? I think it's just an arithmetic error on my part. THanks guys!

OpenStudy (asnaseer):

yw :) - crosses fingers in the hope that he did make a mistake :)

OpenStudy (asnaseer):

*did NOT ...

OpenStudy (anonymous):

Wait, one last thing, when I find the determinants of the 3 x 3, can I just use the diagonal method (I'm not sure of the name) or do I have to use minors?

OpenStudy (asnaseer):

as @TuringTest said, use the method of cofactors - see the links we gave you.

OpenStudy (anonymous):

Ok, once again, thank you :)

OpenStudy (turingtest):

the problem in what you did is that you forgot that all elements on the diagonals will give positive minors.

OpenStudy (turingtest):

I phrased that poorly, not sure how to say it without giving away the answer

OpenStudy (asnaseer):

Also, determinants are usually written with straight lines on either side of the matrix elements - instead of the square brackets that you used.

OpenStudy (anonymous):

Oh! I got it! And @asnaseer for some reason, I see them written both ways when doing school work?

OpenStudy (asnaseer):

Hmmm - ok, well I was taught that straight lines is what should be used. Maybe the notation differs from country to country?

OpenStudy (anonymous):

Perhaps.

OpenStudy (turingtest):

if it's left in brackets it should at least say \(\det\) in front of it In the US we use what @asnaseer said

OpenStudy (asnaseer):

Here is an example of how we are taught in the UK: http://www.intmath.com/matrices-determinants/matrix-determinant-intro.php

OpenStudy (asnaseer):

Also, when working with algebra, if you have a matrix A, then we would write its determinant as either:\[|A|\]or:\[det(A)\]